OpenStudy (anonymous):

The equations for two sides of a parallelogram are -x+10y-42=0 and 3x-y+10=0. Find the equations for the other two sides given that the point (6, -1) is a vertex of the parallelogram, Give your answer in general form.

OpenStudy (phi):

Did you make any progress?

OpenStudy (phi):

I would put both equations into the form y = mx + b so that you can see what the slope of each line is (the "m" value)

OpenStudy (phi):

Maybe make a rough sketch of the lines then plot the point (6,-1) the other two sides of the parallelogram will be parallel to the given lines, and go through the given point (6,-1) you could use the point-slope formula y - y0 = m( x - x0) to find the equation of these two sides. then translate to general form

OpenStudy (phi):

I think general form is Ax + By + C = 0

OpenStudy (anonymous):

@phi OHHH.Thank you again :)